# A minimal model for multigroup adaptive SIS epidemics

### v1

@@ 4.3 Case study 3: Internal and external connections
The constant $c$ allows for the balancing of the importance of internal and external responses to the epidemic. For $c=1$, the link-breaking functional response for external and internal connections is equivalent. Otherwise, for $c>1$, links are broken at a higher rate within communities and for $c<1$, links are broken at a higher rate between communities. The remainder of the parameters as well as the complete graph are equivalent to Case Study 2. Figure 8 depicts the time-evolving prevalence and link densities for three different situations. As expected, high link-breaking rates within communities results in less links within communities, and vice versa.
@@
Figure 8: Case study 3: influence of breaking within or between communities on the prevalence and link densities. The parameters and functional responses are the same as in Case study 2 with a complete network, resulting in $R_{0}=2.53$, with the exception that we used for (a), $c=4$, for (b), $c=1$ and for (c), $c=0.25$.
@@
Figure 9: Case study 3: impact of the parameter $c$ on the steady-state prevalence $\bar{y}_{\infty}$ and the peak prevalence $y_{p}$ in the (a) complete graph and (b) cycle graph. Parameters are the same as Figure 8, except for (b), we have chosen $\beta_{ij}\sim U([0,1])$, such that $R_{0}=1.83$. The horizontal axis is shown on a logarithmic scale.
@@
4.4 Case study 4: Local versus global awareness
@@
As in case study 2 and 3, the link-creation functional responses are taken as
@@
Figure 10: Case study 4: balancing between local and global awareness of the disease (a) within communities and (b) between communities showing the peak prevalence (red triangles) and the steady-state prevalence (green circles). All curves are declining for increasing $c$ values, but the steady-state prevalence within communities is non-monotonic. Simulations are based on an Barabási-Albert graph with $m_{0}=3,m=2$, $\delta_{i}=1,\zeta_{ij}\sim U([0,2]),\xi_{ij}\sim U([0,1])$, $y_{i}(0)=0$ for all $i\neq 1$ and $y_{1}(0)=0.2$ and $z_{ij}(0)=0$ for all $i,j$. Furthermore, (a) $n=20$ nodes, $\beta_{ij}\sim U([0,1])$ such that $R_{0}=2.93$ and (b) $n=50$ nodes, $\beta_{ij}\sim U([0,0.8])$ such that $R_{0}=2.60$.
@@ 5 Conclusions
Lastly, we enhanced our analytical results through an extensive numerical exploration of various case studies. In particular, we showed that the multigroup aNIMFA model allows for periodic orbits in a network with just two communities. We showcased this fact with an asymmetric on/off strategy for the adaptivity between communities, representing total lockdowns when the prevalence reaches a given threshold. Moreover, we investigated the effect of a seasonally-varying infection rate on the prevalence by comparing multiple kinds of networks. Then we showed that breaking connections between communities is preferable over breaking connections within communities in dense graphs, but in sparse networks (with few links) both link-breaking strategies are viable approaches to lessen the disease prevalence. Lastly, we measured the trade-off between local and global awareness on the current state of the epidemic. In many scenarios, using local information appears superior to using global awareness, but further research on larger networks is required to fully support this claim.
@@
Moreover, we remark that our approach to the modelling of adaptivity in the context of epidemic spreading can be easily generalized to more complex models. The main underlying hypothesis of the systems presented here and in [1] is the SIS compartmental structure, which does not allow for any kind of immunity. However, modelling the connectivity based on link densities between between nodes and viral states can be included in any multigroup model, as a level of adaptation of the communities to the evolution of an epidemic. In more complex models, such as the SIR (Susceptible – Infected – Recovered), SIRS, SAIRS (Susceptibel – Asymptomatic – symptomatic Infectious – Recovered – Susceptible) or SIRWS (Susceptible – Infectious – Recovered – Waning immunity – Susceptible), incorporating disease awareness and adaptivity in the system is crucial for improving its realism. Our approach to the subject is very simple, and therefore well-suitable for wide applicability in epidemics.

### v3

@@ 4.2 Case study 2: Internal and external connections
The constant $c$ allows for the balancing of the importance of internal and external responses to the epidemic. For $c=1$, the link-breaking functional response for external and internal connections is equivalent. Otherwise, for $c>1$, links are broken at a higher rate within communities and for $c<1$, links are broken at a higher rate between communities. The remainder of the parameters and the structure of the complete graph can be found in the codes provided in Section 6. Figure 3 depicts the time-evolving prevalence and link densities for three different situations. As expected, high link-breaking rates within communities results in less links within communities, and vice versa.
@@
Figure 3: Case study 2: influence of breaking within or between communities on the prevalence and link densities. For each figure we use $n=20$ and the values of parameters were obtained by sampling, respectively: $\beta_{ij}\sim U([0,0.25]),\delta_{i}=1,\zeta_{ij}\sim U([0,2]),\xi_{ij}\sim U([0,1])$ for all $i,j$, where $U$ denotes the uniform distribution, resulting in $R_{0}=2.53$, with the exception that we used for (a), $c=4$, for (b), $c=1$ and for (c), $c=0.25$.
@@
Figure 4: Case study 2: impact of the parameter $c$ on the steady-state prevalence $\bar{y}_{\infty}$ and the peak prevalence $y_{p}$ in the (a) complete graph and (b) cycle graph. Parameters are the same as Figure 3, except for (b), we have chosen a different sampling for $\beta_{ij}\sim U([0,1])$, where $U$ denotes the uniform distribution, such that $R_{0}=1.83$. The horizontal axis is shown on a logarithmic scale.
@@
4.3 Case study 3: Local versus global awareness
@@
As in case study 2, the link-creation functional responses are taken as
@@
Figure 5: Case study 3: balancing between local and global awareness of the disease (a) within communities and (b) between communities showing the peak prevalence (red triangles) and the steady-state prevalence (green circles). All curves are declining for increasing $c$ values, but the steady-state prevalence within communities is non-monotonic. Simulations are based on an Barabási-Albert graph with $m_{0}=3,m=2$, with the values of parameters sampling: $\delta_{i}=1,\zeta_{ij}\sim U([0,2]),\xi_{ij}\sim U([0,1])$, $y_{i}(0)=0$ for all $i\neq 1$ and $y_{1}(0)=0.2$ and $z_{ij}(0)=0$ for all $i,j$, where $U$ denotes the uniform distribution. Furthermore, (a) $n=20$ nodes, $\beta_{ij}\sim U([0,1])$ such that $R_{0}=2.93$ and (b) $n=50$ nodes, $\beta_{ij}\sim U([0,0.8])$ such that $R_{0}=2.60$.
@@ 5 Conclusions
Lastly, we enhanced our analytical results through an extensive numerical exploration of various case studies. In particular, we showed that the multigroup aNIMFA model allows for periodic orbits in a network with just two communities. We showcased this fact with an asymmetric on/off strategy for the adaptivity between communities, representing total lockdowns when the prevalence reaches a given threshold. Then, we showed that breaking connections between communities is preferable over breaking connections within communities in dense graphs, but in sparse networks (with few links) both link-breaking strategies are viable approaches to lessen the disease prevalence. Lastly, we measured the trade-off between local and global awareness on the current state of the epidemic. In many scenarios, using local information appears superior to using global awareness, but further research on larger networks is required to fully support this claim.
@@
Moreover, we remark that our approach to the modelling of adaptivity in the context of epidemic spreading can be easily generalized to more complex models. The main underlying hypothesis of the systems presented here and in [1] is the SIS compartmental structure, which does not allow for any kind of immunity. However, modelling the connectivity based on link densities between between nodes and viral states can be included in any multigroup model, as a level of adaptation of the communities to the evolution of an epidemic. In more complex models, such as the SIR (Susceptible – Infected – Recovered), SIRS, SAIRS (Susceptible – Asymptomatic – symptomatic Infectious – Recovered – Susceptible) or SIRWS (Susceptible – Infectious – Recovered – Waning immunity – Susceptible), incorporating disease awareness and adaptivity in the system is crucial for improving its realism. Our approach to the subject is very simple, and therefore well-suitable for wide applicability in epidemics. In particular, one could consider a variation of system (2) in which some of the parameters, e.g. $\beta_{ij}$, vary periodic in time, mimicking seasonality.
Lastly, one may consider the inclusion of time-delayed mechanisms and multiple viruses spreading in the same population. For the former, it could be interesting to investigate an adaptivity based on e.g. yesterday’s values on the prevalence, which are more realistically available than real-time ones. For the latter, one could have different adaptivity strategies depending on how each virus is perceived by the population, greatly increasing the complexity of the dynamics.
6 Data Availability Statement
All the codes used to generate Figures 2-5 are avaiable in the GitHub repository
https://github.com/SaraSottile/AdaptativeMeanField.